Find the 50th term of this arithmetic sequence:

2, 5, 8, 11, . . .
1
Expert's answer
2019-05-28T03:40:41-0400

The first term is 2.

The second term is 5 = 2 + 3.

The third term is 8 = 5 + 3 = 2 + 3 * 2, etc.

Then the n-th term will be 2 + 3(n-1).

So the 50-th member will be equal to 2+3*(50-1)=2+3*49=149.